# Exploring Holomorphic Retracts

### v1

@@ Abstract.
The purpose of this article is towards systematically characterizing holomorphic retracts of domains of holomorphy; to begin with, bounded balanced pseudoconvex domains $B\subset\mathbb{C}^{N}$. Specifically, we show that every (holomorphic) retract of $B$ passing through its center (i.e., the origin), is the graph of a holomorphic map over a linear subspace of $B$. If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics; the necessity of the boundary points to be extreme points is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well. This revisit of Vesentini’s result leads to a small modification that is useful to provide interesting examples. As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ which is submersive at its center, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$. To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of strictly convex balls, with respect to various norms and obtain a complete description of retracts passing through the origin, generalizing that of the polydisc due to Heath and Suffrdige and giving a simpler proof as well. The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$. Along the way, we discuss examples of non-convex but pseudoconvex bounded balanced domains in which most one-dimensional linear subspaces fail to be holomorphic retracts, in good contrast to the convex case. Indeed, for the simplest possible case here namely the ‘$\ell^{q}$-ball’ for $10<q<$ 1, we obtain a complete list of all retracts through the origin; this enables an illustration of applying retracts to establishing biholomorphic inequivalences. To go beyond balanced domains, we then first obtain a complete characterization of retracts of the Hartogs triangle and ‘analytic complements’ thereof. Thereafter, we drop the assumption of boundedness as well, to attain similar characterization results for domains which are neither bounded nor topologically trivial. We conclude by reporting some results on the retracts of $\mathbb{C}^{2}$.
@@ 1. Introduction
We begin with remarks for motivation and background, possibly a bit elaborate; this owes to the unavailability of a survey or expository article on holomorphic retracts to refer the reader to. This will also simultaneously help us set up notations and fix terminologies. By a retraction we shall mean an idempotent map, whereas by a retract, we mean the image of such a map. More precisely, we shall be interested in holomorphic retracts i.e., the image of a holomorphic retraction which by definition, is a holomorphic self-map $\rho$ of a domain $D\subset\mathbb{C}^{N}$ (or more generally, a complex manifold) satisfying the idempotent condition $\rho^{\circ 2}:=\rho\circ\rho=\rho$. This means that $\rho$ acts as the identity map on $Z=\rho(D)$ which shall be our notation for a retract throughout this article; a trivia worth noting is that $Z$ is precisely the set of points of $D$ fixed by $\rho$. A pair of extreme or trivial examples of retractions (resp. retracts) of any complex manifold are the identity map and constant maps (resp. $D$ and its points); obviously, we would not like to talk of them any further and thereby shall be referred to as trivial retracts henceforth. For instance, with this terminology and understanding, we shall say that in dimension $N=1$, there are no non-trivial retracts. This means that for any domain $D$ in $\mathbb{C}$ or any Riemann surface, any holomorphic self-map $\rho$ satisfying $\rho^{\circ 2}=\rho$ is either a constant map or else, the identity map on $D$. Indeed, this follows from the open mapping theorem together with, the idempotency condition which implies that $\rho(D)$ is a closed subset of $D$. In other words, holomorphic retracts are interesting only when the dimension $N$ gets bigger than one. Now in multidimensional complex analysis, a lot of effort has gone into the study of biholomorphic and more generally, proper (holomorphic [^1] Unless otherwise mentioned, all maps in this article are to be understood to be holomorphic.) maps, as is natural. To mention just one or two of the first basic facts about proper maps between a pair of domains in $\mathbb{C}^{N}$, they are known to be always surjective and for many a class of domains $D$ with ‘nice’ boundaries, proper self-maps turn out to be automorphisms of $D$. Though absolutely trivial to check, let us mention for contrast with the just mentioned facts about proper maps, that: non-trivial retractions are never surjective (and never are they, automorphisms!); as one another sharp contrast, proper maps never decrease dimension, whereas non-trivial retractions always decrease dimension. In other words, retractions could be viewed as being, loosely speaking, at the ‘opposite end of the spectrum’ when compared to proper maps.
@@
While much work has been carried out on proper holomorphic maps and their images, far less has investigation has been done about retracts. Infact, in addition to (the foregoing remarks and) their role of in certain fundamental mapping problems above, retracts are of significance in complex dynamics as well. To just be specific by providing a standard result in this regard, we mention the following theorem which shows how retracts arise in the study of iterates of a holomorphic self-map of taut manifolds. Needless to say, this theorem has been applied to the study of the special case of proper holomorphic self-maps. In light of the the stark contrasts between proper maps and retraction maps mentioned above, the connection amongst these maps, is thereby all the more interesting.
Theorem 1.1 (cf. Bedford, [Bdfrd]; Abate, [Abate_Horosphre]).
@@
Let us then begin with the obvious observation that retracts are always fixed point sets of some holomorphic map; but the converse need not be true in general. But then again, an easy but noteworthy consequence of theorem 1.1 is that the fixed point set of every holomorphic self-map of any taut manifold $X$, is contained within a non-trivial retract. Furthermore, there are results for an exact equality rather than just containment, if so sought for: as shown in [Mazet_Vigue_fxdpnt], [Thai_Fxdpnt_cnvx] among others, the fixed point set of any holomorphic self map of any hyperbolic convex domain in $\mathbb{C}^{N}$ – whenever it is non-empty – is also a retract of the domain. Whether or not fixed point sets of a holomorphic self map is the retract of possibly some other holomorphic self map of $X$, a definitive basic feature that all retracts have in common is that every retract $Z$ has the extension property in the sense that every holomorphic function on $Z$ admits (just by composition with the retraction mapping $X$ onto $Z$) a holomorphic extension to $X$. When the given function on $Z$ is bounded, we also have (by the same means of composition with the retraction map), a holomorphic extension on $X$ with the same sup-norm and thus retracts form the simplest class of subvarieties of $X$ with the norm-preserving extension property to $X$. Whether retracts are the only subvarieties with this property was examined and shown to be so, in many special cases by Kosiński and McCarthy in the relatively recent work [Kosnski] wherein counterexamples have also been given. All this is only towards indicating the significance and motivation for the study of holomorphic retracts, in general. Our first goal for this article is rather specific and it is: to obtain a description of the holomorphic retracts of pseudoconvex bounded balanced domains passing through its center, namely the origin in $\mathbb{C}^{N}$. This itself is noteworthy; if not for anything else, this suffices for an important application to attain (employing the above theorem 1.1), the fairly recent result of [JJ] together with [Pranav_JJ] that: every proper holomorphic self-map of any smoothly bounded balanced pseudoconvex domain of finite type is an automorphism. In-fact our methods help simplify a part of the proof of this result which we shall not however elaborate here.
@@
As mentioned in the abstract, we intend a systematic presentation; so, to not be amiss about addressing the simplest of questions in the aforementioned first goal then, we recall that in [Vesntn], E. Vesentini already proved the following fundamental result about the form of the retracts of certain convex domains. More important than the various hypotheses of this result, as pointed out by J. P. Vigue in [Vigue_fxdpnt_cnvx], Vesentini’s approach makes ‘heavy’ use of complex geodesics. We shall to begin with, give a relatively elementary proof of this result without using complex geodesics. It may be noted that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal and thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$ but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal while it is indeed $\mathbb{C}$-extremal. The notions of $\mathbb{R}$-extremality, $\mathbb{C}$-extremality and the relationship between them is recalled in the following section 2 about preliminaries.
@@
Now, the condition of convexity on $D$, which entails a certain condition to be satisfied by the entire boundary $\partial D$, can be dropped in various ways. To begin with here, let us only remark that we can do so in the following manner, which will actually be useful in dealing with certain concrete (counter-)examples in the sequel. Specifically, we shall show the following (without using ‘complex geodesics’).
@@
(A small modification of Vesentini’s result): Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a retract through the origin with the property that $D\cap T_{0}Z$ is a convex domain in $T_{0}Z$ with $\mathbb{C}$-extremal boundary, then $Z$ is a linear subspace of $D$. In particular, every $1$-dimensional retract through the origin of a bounded balanced pseudoconvex domain $D$ with $\mathbb{C}$-extremal boundary, is linear.
We emphasize that this result is not really new and our inclusion of this is towards a relatively self-contained exposition, and more importantly, in the proof of proposition 5.1 below. Indeed, we would not like to be amiss to observe associated facts which are not explicitly recorded in the literature but can be attained by reading proofs of results carefully (and such observations are actually useful for us much later in the sequel): for instance, the apparently strong condition (when $\dim(Z)\geq 2$) on the convexity of $D\cap T_{0}Z$ in the above theorem can be dropped while attaining the same conclusion; specifically, we have the following.
@@
(A small modification of proposition $4.5$ of [JJ]). Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a retract through the origin with the property that the boundary points of $D:=D\cap T_{0}Z$ viewed as a domain in $T_{0}Z$ are holomorphically extreme. Then, $Z$ is linear.
Amongst other results about retracts, this helps in establishing biholomorphic inequivalences, particularly with product domains. Stated differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., biholomorphic to a product domain. Let us only indicate the key idea here namely, the fundamental common feature amongst all product domains that: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed. Let us just state precisely here the simplest result to this effect to indicate what this means; leaving more general formulations to be laid out in section 7 wherein we shall have other remarks to tie up with.
@@
An extension of this to higher dimensions is briefly discussed in section 7. We also hasten to mention that there are better techniques to detect the occurrence of a product structure for domains far beyond the class of balanced domains owing particularly to the very recent work [Bh-Bo-Su], which however requires some quantitative knowledge about the squeezing function. However we believe that for balanced domains, the above proposition gives a much simpler approach in practice, since a balanced domain is most often given by its Minkowski functional. Needless to say, there are plenty of balanced domains of holomorphy satisfying the non-convexity condition in the above proposition, as will also be discussed in section 7.
@@
Getting back to retracts, let us start by noting that while Vesentini’s theorem says that all retracts of $D$ (as in that theorem) passing through the origin, are linear, it does not say that all linear subspaces are retracts. Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]). Indeed, the question of precisely which linear subspaces are realizable as retracts of $D$ as in the above theorem, is quite an interesting question and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections (when we measure the norm of the projection with respect to the operator norm induced by $\mu$)? The answer turns out to vary even among the $l^{p}$ norms i.e., depend upon $p$ when we take $\mu=\|\cdot\|_{l^{p}}$ even if we exclude the case $p=2$, which may not be surprising at all for this case is distinguished as the one that arises from an inner product; specifically, the case $p=\infty$ is also different from the other cases. Before getting to such differences, let’s mention a unified result. As our interest lies in retracts, we shall phrase in it in the language of retracts rather than norm-one projections. Specifically, the characterization of linear retracts of the standard $l^{p}$ ball ($1\leq p<\infty,~{}p\neq 2$) can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
@@
(theorem 2.a.4 of [Lindstraus]): Let $Y$ be a linear subspace of $\mathbb{C}^{N}$ equipped with $\|\cdot\|_{l^{p}}$ where $1\leq p<\infty,p\neq 2$. Then $Y\cap B_{\ell^{p}}$ is a linear retract of $B_{\ell^{p}}$ iff there exist vectors $\{u_{j}\}_{j=1}^{m}$ of norm 1 in $\mathbb{C}^{N}$ of the form
@@
such that $Y=span\{u_{j}\}_{j=1}^{m}$, where $\sigma_{j}:=supp(u_{j})=\{1\leq i\leq N:u_{j}(i)\neq 0\}$ (and $\lambda_{i}$’s are some complex numbers).
The case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, in which all linear subspaces are retracts. We would rather like to draw attention to the more important fact that this result is not true for $p=\infty$ to demonstrate which (by a counterexample as simple as possible), we include an example in section 3. This naturally leads to the question as to whether it is possible at all to pin down precisely which linear subspaces of the polydisc $\Delta^{N}$ can be realized as linear retracts (recall that the polydisc is the exceptional case corresponding to $p=\infty$ which is left out by theorem 1.7). Towards this, we first note the following characterization of those linear subspaces of $\Delta^{N}$ which are its retracts. Indeed, this can readily derived from [Heath_Suff], though it is not stated in these explicit terms; contained within our treatment will be a simpler proof of the main result of [Heath_Suff] about the characterization of all retracts of the polydisc as well as a deduction of the following special one about its linear retracts.

### v6

@@ Abstract.
The purpose of this article is towards systematically characterizing holomorphic retracts of domains of holomorphy; to begin with, bounded balanced pseudoconvex domains $B\subset\mathbb{C}^{N}$. Specifically, we show that every (holomorphic) retract of $B$ passing through its center (i.e., the origin), is the graph of a holomorphic map over a linear subspace of $B$. If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics; the necessity of the boundary points to be extreme points is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well. This revisit of Vesentini’s result leads to a small modification that is useful to provide interesting examples; specifically, in completely determining the retracts through the origin of certain balanced domains with analytic varieties in their boundaries. The converse of Vesentini’s theorem is established and general results characterizing the non-linear retracts through the origin are attained (for fairly general domains in this context) when the boundary fails to have extreme points. As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ which is submersive at its center, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$. To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of strictly convex balls, with respect to various norms and obtain a complete description of retracts passing through the origin, generalizing that of the polydisc due to Heath and Suffridge and giving a simpler proof as well. The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$. Along the way, we discuss examples of non-convex but pseudoconvex bounded balanced domains in which most one-dimensional linear subspaces fail to be holomorphic retracts, in good contrast to the convex case. Indeed, for the simplest possible case here namely the ‘$\ell^{q}$-ball’ for $10<q<$ 1, we obtain a complete list of all retracts through the origin of each possible dimension; this enables an illustration of applying retracts to establishing biholomorphic inequivalences. To go beyond balanced domains, we then first obtain a complete characterization of retracts of the Hartogs triangle and ‘analytic complements’ thereof. Thereafter, we drop the assumption of boundedness as well, to attain similar characterization results for domains which are neither bounded nor topologically trivial. We conclude with some expositions about retracts of $\mathbb{C}^{2}$.
@@ 1. Introduction
We begin with remarks for motivation and background, possibly a bit elaborate; this owes to the unavailability of a survey or expository article on holomorphic retracts to refer the reader to. This will also simultaneously help us set up notations and fix terminologies. We emphasize that this article is a survey in parts, striving to give references for the known results (to the best of our knowledge). By a retraction we shall mean an idempotent map, whereas by a retract, we mean the image of such a map. More precisely, we shall be interested in holomorphic retracts i.e., the image of a holomorphic retraction which by definition, is a holomorphic self-map $\rho$ of a domain $D\subset\mathbb{C}^{N}$ (or more generally, a complex manifold) satisfying the idempotent condition $\rho^{\circ 2}:=\rho\circ\rho=\rho$. This means that $\rho$ acts as the identity map on $Z=\rho(D)$ which shall be our notation for a retract throughout this article; a trivia worth noting is that $Z$ is precisely the set of points of $D$ fixed by $\rho$. A pair of extreme or trivial examples of retractions (resp. retracts) of any complex manifold are the identity map and constant maps (resp. $D$ and its points); obviously, we would not like to talk of them any further and thereby shall be referred to as trivial retracts henceforth. For instance, with this terminology and understanding, we shall say that in dimension $N=1$, there are no non-trivial retracts. This means that for any domain $D$ in $\mathbb{C}$ or any Riemann surface, any holomorphic self-map $\rho$ satisfying $\rho^{\circ 2}=\rho$ is either a constant map or else, the identity map on $D$. Indeed, this follows from the open mapping theorem together with, the idempotency condition which implies that $\rho(D)$ is a closed subset of $D$. In other words, holomorphic retracts are interesting only when the dimension $N$ gets bigger than one. Now in multidimensional complex analysis, a lot of effort has gone into the study of biholomorphic and more generally, proper (holomorphic [^1] Unless otherwise mentioned, all maps in this article are to be understood to be holomorphic.) maps, as is natural. To mention just one or two of the first basic facts about proper maps between a pair of domains in $\mathbb{C}^{N}$, they are known to be always surjective and for many a class of domains $D$ with ‘nice’ boundaries, proper self-maps turn out to be automorphisms of $D$. Though absolutely trivial to check, let us mention for contrast with the just mentioned facts about proper maps, that: non-trivial retractions are never surjective (and never are they, automorphisms!); as one another sharp contrast, proper maps never decrease dimension, whereas non-trivial retractions always decrease dimension. In other words, retractions could be viewed as being, loosely speaking, at the ‘opposite end of the spectrum’ when compared to proper maps.
@@
While much work has been carried out on proper holomorphic maps and their images, far less investigation has been done about retracts. Infact, in addition to (the foregoing remarks and) their role of in certain fundamental mapping problems above, retracts are of significance in complex dynamics as well. To just be specific by providing a standard result in this regard, we mention the following theorem which shows how retracts arise in the study of iterates of a holomorphic self-map of taut manifolds. Needless to say, this theorem has been applied to the study of the special case of proper holomorphic self-maps. In light of the the stark contrasts between proper maps and retraction maps mentioned above, the connection amongst these maps, is thereby all the more interesting.
Theorem 1.1 (Bedford, [Bdfrd]; Abate, [Abate_Horosphre]).
@@
Let us then begin with the obvious observation that retracts are always fixed point sets of some holomorphic map; but the converse need not be true in general. But then again, an easy but noteworthy consequence of theorem 1.1 is that the fixed point set of every holomorphic self-map of any taut manifold $X$, is contained within a non-trivial retract. Furthermore, there are results for an exact equality rather than just containment, if so sought for: as shown in [Mazet_Vigue_fxdpnt], [Thai_Fxdpnt_cnvx] among others, the fixed point set of any holomorphic self map of any hyperbolic convex domain in $\mathbb{C}^{N}$ – whenever it is non-empty – is also a retract of the domain. Whether or not fixed point sets of a holomorphic self map is the retract of possibly some other holomorphic self map of $X$, a definitive basic feature that all retracts have in common is that every retract $Z$ has the extension property in the sense that every holomorphic function on $Z$ admits (just by composition with the retraction mapping $X$ onto $Z$) a holomorphic extension to $X$. When the given function on $Z$ is bounded, we also have (by the same means of composition with the retraction map), a holomorphic extension on $X$ with the same sup-norm and thus retracts form the simplest class of subvarieties of $X$ with the norm-preserving extension property to $X$. Whether retracts are the only subvarieties with this property was examined and shown to be so, in many special cases by Kosiński and McCarthy in the relatively recent work [Kosnski] wherein counterexamples have also been given. All this is only towards indicating the significance and motivation for the study of holomorphic retracts, in general. Our first goal for this article is rather specific and it is: to obtain a description of the holomorphic retracts of pseudoconvex bounded balanced domains passing through its center, namely the origin in $\mathbb{C}^{N}$. This itself is noteworthy; if not for anything else, this suffices for an important application to attain (employing the above theorem 1.1), the fairly recent result of [JJ] together with [Pranav_JJ] that: every proper holomorphic self-map of any smoothly bounded balanced pseudoconvex domain of finite type is an automorphism. We shall not however elaborate upon this here.
@@
As mentioned in the abstract, we intend a systematic presentation; so, to not be amiss about addressing the simplest of questions in the aforementioned first goal then, we recall that in [Vesntn], E. Vesentini already proved the following fundamental result about the form of the retracts of certain convex domains. More important than the various hypotheses of this result, as pointed out by J. P. Vigue in [Vigue_fxdpnt_cnvx], Vesentini’s approach makes ‘heavy’ use of complex geodesics. We shall to begin with, give a relatively elementary proof of this result without using complex geodesics. It may be noted that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal; thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$, but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal but indeed $\mathbb{C}$-extremal. The notions of $\mathbb{R}$-extremality, $\mathbb{C}$-extremality and the relationship between them is recalled in the following section 2 about preliminaries. The necessity of such extremality hypotheses is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well.
@@
Now, the condition of convexity on $D$, which entails a certain condition to be satisfied by the entire boundary $\partial D$, can be dropped in various ways. To begin with here, let us only remark that we can do so in the following manner; this will actually be useful in dealing with certain concrete examples in the sequel (example 8.7 in section 8). This is one another major reason for in revisiting this result of Vesentini. Specifically, we shall show the following without using complex geodesics as alluded to earlier.
@@
(A small modification of Vesentini’s result): Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a retract through the origin with the property that: $D\cap T_{0}Z$ is convex with points of $\partial D\cap T_{0}Z$ being $\mathbb{C}$-extremal boundary points of $D$; then $Z$ is a linear subspace of $D$. In particular, every $1$-dimensional retract through the origin of a bounded balanced pseudoconvex domain $D$ with $\mathbb{C}$-extremal boundary, is linear.
Our inclusion of this result first towards a relatively self-contained exposition, has other more important motivations than just a more elementary proof: to furnish examples of balanced non-convex (but pseudoconvex) domains possessing non-trivial analytic varieties in the boundary, wherein the retracts through the center can be completely determined. This is accomplished in section 8 wherein the above modified version of Vesentini’s result has also a role to play in other propositions there. Also in general, we would not like to be amiss to observe associated facts which are not explicitly recorded in the literature but can be attained by reading proofs of results carefully – these are mostly exposited in the ‘preliminaries’ section; more importantly, such observations are actually useful for us much later in the sequel. While we relegate such observations to later, one that we think must be mentioned right away here, is the affirmative answer to the question about the very existence of atleast one retract through the origin of an arbitrary bounded balanced domain of holomorphy without any further assumptions. This is verified in proposition 2.35. Getting back to the above theorem let us now mention for instance, the apparently strong condition (when $\dim(Z)\geq 2$) on the convexity of $D\cap T_{0}Z$ in the above theorem can be dropped while attaining the same conclusion; specifically, we have the following.
@@
(A small modification of proposition $4.5$ of [JJ]): Let $D$ be any balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $D\cap T_{0}Z$ is bounded and $Z$ is a retract through the origin with the property that all points of $\partial D\cap T_{0}Z$ are holomorphically extreme boundary points of $D$. Then, $Z$ is linear.
Given this, it may seem questionable as to why we have stated theorem 1.3 in this introductory section; in this regard let us highlight the existence of domains as already alluded to, whose boundaries are not holomorphically extreme but are $\mathbb{C}$-extremal. Also, $\mathbb{C}$-extremality is far easier to verify than holomorphic extremality which apparently is a more restrictive condition. Indeed, a concrete example of a domain of the kind just mentioned is recorded in proposition 8.8 wherein the domain also serves to show that the converse of the above proposition is false. As for the significance of the above proposition 1.4 itself, let us only remark that it will be employed in determining retracts through the origin of all possible dimensions in $\ell_{q}$-‘balls’ as in theorem 1.9 below. It also helps in establishing biholomorphic inequivalences, particularly with product domains; for instance, a domain as in the above proposition with holomorphically extreme boundary and the polydisc – this is only to give a quick example here. To get new results, we must consider domains with greater complexity, to proceed systematically towards which we deal first with a simple class of domains which allow analytic varieties in their boundaries. Stating the essence here differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., biholomorphic to a product domain. Let us only indicate the key idea here namely, the fundamental common feature amongst all product domains that: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed. Let us just state precisely here the simplest result to this effect to indicate what this means; leaving more general formulations to be laid out in section 8 wherein we shall have other remarks to tie up with.
@@
Domains as in the above proposition may be regarded as being maximally non-convex in the realm of bounded balanced domains, remembering that it is impossible to have a bounded domain which is non-convex throughout the boundary i.e., at every boundary point. While convex domains is a much studied class, the other extreme has by far less been; this proposition contributes towards that. An extension of this to higher dimensions is briefly discussed in section 8. We also hasten to mention that there are better techniques to detect the occurrence of a product structure for domains far beyond the class of balanced domains owing particularly to the very recent work [Bh-Bo-Su], which however requires some quantitative knowledge about the squeezing function. However we believe that for balanced domains, the above proposition gives a much simpler approach in practice, since a balanced domain is most often given by its Minkowski functional. Needless to say, there are plenty of balanced domains of holomorphy satisfying the non-convexity condition in the above proposition, as will also be discussed in section 8.
@@
Getting back to retracts, let us start by noting that while Vesentini’s theorem says that all retracts of $D$ (as in that theorem) passing through the origin, are linear, it does not say that all linear subspaces are retracts. Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]). The question of precisely which linear subspaces are realizable as retracts of $D$ as in the above theorem is significant for us. For instance: while every retract of the polydisc $\Delta^{N}$ is a graph, not every graph in $\Delta^{N}$ is a retract; only those which are graphs on linear retracts of $\Delta^{N}$ are, making it important to precisely pin down which linear subspaces are retracts. Indeed, this is quite an interesting question in its own right and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections (when we measure the norm of the projection with respect to the operator norm induced by $\mu$)? The answer turns out to vary even among the $l^{p}$ norms i.e., depend upon $p$ when we take $\mu=\|\cdot\|_{l^{p}}$ even if we exclude the case $p=2$, which may not be surprising at all for this case is distinguished as the one that arises from an inner product; specifically, the case $p=\infty$ is also different from the other cases. Before getting to such differences, let’s mention a unified result. As our interest lies in retracts, we shall phrase in it in the language of retracts rather than norm-one projections. Specifically, the characterization of linear (thereby holomorphic, owing to Vesentini’s theorem) retracts of the standard $l^{p}$ ball ($1\leq p<\infty,~{}p\neq 2$) can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
@@
(Theorem 2.a.4 of [Lindstraus]): Let $Y$ be a linear subspace of $\mathbb{C}^{N}$ equipped with $\|\cdot\|_{l^{p}}$ where $1\leq p<\infty,p\neq 2$. Then $Y\cap B_{\ell^{p}}$ is a linear retract of $B_{\ell^{p}}$ if and only if there exist vectors $\{u_{j}\}_{j=1}^{m}$ of $\ell^{p}$-norm 1 in $\mathbb{C}^{N}$ of the form
@@
such that $Y=span\{u_{j}\}_{j=1}^{m}$, where $\sigma_{j}:=supp(u_{j})=\{1\leq i\leq N:u_{j}(i)\neq 0\}$ (and $\lambda_{i}$’s are some complex numbers such that the $\ell^{p}$-norm of $u_{j}$ are equal to $1$). Here $e_{i}$ denotes the standard basis of $\mathbb{C}^{N}$.
We shall discuss alternative formulations of the above result which are more geometrically revealing in section 5. In-particular, retracts of $\ell^{p}$-balls are isometric copies of lower dimensional standard $\ell^{p}$-balls. Next, we note that the case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, in which all linear subspaces are retracts. We would rather like to draw attention to the more important fact that this result is not true for $p=\infty$ to demonstrate which (by a counterexample as simple as possible), we include an example in section 3. This naturally leads to the question as to whether it is possible at all to pin down precisely which linear subspaces of the polydisc $\Delta^{N}$ can be realized as linear retracts (recall that the polydisc is the exceptional case corresponding to $p=\infty$ which is left out by theorem 1.7). Towards this, we first note the following characterization of those linear subspaces of $\Delta^{N}$ which are its retracts. Indeed, this can readily derived from [Heath_Suff], though it is not stated in these explicit terms; contained within our treatment will be a simpler proof of the main result of [Heath_Suff] about the characterization of all retracts of the polydisc as well as a deduction of the following special one about its linear retracts.